NNCuts on LRHWW, Mll and MetSig

Richard St. Denis University of Glasgow/CDF 17 December 2007

H → WW Meeting

[figure]

Outline

▶ MetSig, Met and SumEt by jet and Hi/Lo S/B ▶ BaseSameSign

H → WW Meeting

[figure: SumEt]

Richard St. Denis, Glasgow University/CDF, 10 December 2

H → WW Meeting

[figure: SumEt]

Richard St. Denis, Glasgow University/CDF, 10 December 3

H → WW Meeting

[figure: DiMass]

Richard St. Denis, Glasgow University/CDF, 10 December 4

H → WW Meeting

[figure: DiMass]

Richard St. Denis, Glasgow University/CDF, 10 December 5

H → WW Meeting

[figure: Met]

Richard St. Denis, Glasgow University/CDF, 10 December 6

H → WW Meeting

[figure: Met]

Richard St. Denis, Glasgow University/CDF, 10 December 7

H → WW Meeting

[figure: MetSig]

Richard St. Denis, Glasgow University/CDF, 10 December 8

H → WW Meeting

[figure: MetSig]

Richard St. Denis, Glasgow University/CDF, 10 December 9

H → WW Meeting

[figure]

dPhiLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 10

H → WW Meeting

[figure]

dPhiLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 11

H → WW Meeting

[figure]

dRLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 12

H → WW Meeting

[figure]

dRLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 13

H → WW Meeting

[figure: jet1 Et]

Richard St. Denis, Glasgow University/CDF, 10 December 14

H → WW Meeting

[figure: jet1 Et]

Richard St. Denis, Glasgow University/CDF, 10 December 15

H → WW Meeting

[figure]

Base SameSign

Richard St. Denis, Glasgow University/CDF, 10 December 16

H → WW Meeting

[figure]

Base SameSign

Richard St. Denis, Glasgow University/CDF, 10 December 17

H → WW Meeting

[figure: Base]

Richard St. Denis, Glasgow University/CDF, 10 December 18

H → WW Meeting

[figure: Base]

Richard St. Denis, Glasgow University/CDF, 10 December 19

H → WW Meeting

[figure: Base]

Richard St. Denis, Glasgow University/CDF, 10 December 20

H → WW Meeting

[figure: Base]

Richard St. Denis, Glasgow University/CDF, 10 December 21

H → WW Meeting

[figure]

Calibrating SumEt

Richard St. Denis, Glasgow University/CDF, 10 December 22

[figure]

Calibrating SumEt: DY

▶ DY: for CM or No ISR and no jet:

L_{1}+ L_{1} = 0

▶ DY: for ISR and no jet

L′_{1}+ L′_{2} = L_{tot}

▶ DY: for ISR and 1 jet

L′_{1}+ L′_{2} = L_{tot}= ∑ E_{T}+ E_{t}^{jet}

This estabilishes the relationship between observed ∑ E_{T} and the vector to add to the CM vector. If a jet is present, then the table of Jet and ∑ E_{T} needs to be applied. The quantities are ∑ E_{T}^{DY 0}and (∑ E_{T}^{DY 1}, E_{t}^{jet1}).

[figure]

Applying SumEt Calibration:WW

▶ WW: for CM or No ISR and no jet:

L_{1}+ L_{2} = L_{tot}= E/_{T}= E/_{T}_{obs} = −(ν_{1}+ ν_{2})

Here, L_{tot}≡ L_{CM}

▶ WW: for ISR and no jet

L′_{1}+ L′_{2} = L′_{tot}= ∑ E_{T}^{W W}+ L_{tot}

Hence

L_{tot}= L′_{tot}− ∑ E_{T}^{W W}

We assume now that ∑ E_{T}^{W W}= ∑ E_{T}^{DY} For a given ∑ E_{T}^{W W} and substitute in the corresponding calibration for ∑ E_{T}^{DY}.

[figure]

▶ WW: for ISR and 1 jet

L′_{1}+ L′_{2} = L′_{tot}= ∑ E_{T}^{W W}+ E_{t}^{jet} + L_{tot}

Hence

L_{tot}= L′_{tot}− ∑ E_{T}^{W W}− E_{t}^{jet}

We assume now that for (E_{t}^{jet})^{W W}(∑ E_{T}^{W W})_{1jet} = (∑ E_{T}^{DY})_{1jet} For a given (SumEt^{W W})_{1jet} and (E_{t}^{jet})^{W W}substitute in the corresponding calibration for (∑ E_{T}^{DY})_{1jet} and (E_{t}^{jet})^{DY}